# Safe Planning in Unknown Environments using Conformalized Semantic Maps

### v1

@@ III-B Planning over Conformalized Semantic Maps
Consider a mission scenario $\xi=({\mathbf{x}}_{0},\Omega,\psi)$ drawn from $\mathcal{D}$. Our goal is to compute a plan ${\mathbf{x}}_{0:H}$ satisfying $\psi$ with probability at least $1-\alpha$. This path will be constructed incrementally as the map ${\mathbf{m}}_{t}$ is updated online; see Alg. 2. First, we collect the calibration dataset $\{\xi_{i}\}_{i=1}^{D}$, and compute $\hat{s}$ as described in Section III-A [lines 3-4, Alg. 2]. Assume that at time $t>0$, the robot is at state ${\mathbf{x}}_{t}$, collects measurements ${\mathcal{Z}}_{t}$, and constructs the map ${\mathbf{m}}_{t}$ [lines 6-7, Alg. 2]. Using ${\mathbf{m}}_{t}$, the prediction sets ${\mathcal{C}}_{j}({\mathbf{m}}_{t})$ are constructed for all cells $j\in{\mathcal{J}}_{t}$ [line 8, Alg. 2]. We now discuss how the next robot state ${\mathbf{x}}_{t+1}$ is computed using these sets [lines 914, Alg. 2]. The next state ${\mathbf{x}}_{t+1}$ must be chosen so that it satisfies (1). Since the regions ${\mathcal{L}}_{k}$ are unknown, satisfaction of (1) must be checked using the map ${\mathbf{m}}_{t}$. Specifically, given ${\mathbf{m}}_{t}$, (1) can be re-written as:
$\lVert{\mathbf{x}}_{t+1}|\Omega-c_{j}\rVert\geq d_{k_{j}}+r_{m},$ (6)
@@
Exploitation Phase: Using an existing motion planner [1, 2, 3], we first check whether there exists a robot path ${\mathbf{p}}={\mathbf{p}}_{0},{\mathbf{p}}_{1},\dots,{\mathbf{p}}_{T}$, with ${\mathbf{p}}_{n}\in{\mathcal{X}}$ for all $n\in\{0,\dots,T\}$ and some $T\geq 0$ that completes $\psi$ under the assigned labels $k_{j}$ [line 9, Alg. 2]. The path must satisfy ${\mathbf{p}}_{0}={\mathbf{x}}_{t}$, ${\mathbf{p}}_{T}\in{\mathcal{X}}_{g}$, and $\lVert{\mathbf{p}}_{n}|_{\Omega}-c_{j}\rVert\geq d_{k_{j}}+r_{m}$ for all $n$. If such a feasible plan exists, the robot selects ${\mathbf{x}}_{t+1}={\mathbf{p}}_{1}$ and moves towards it, updates the time step to $t+1$, and repeats the process [lines 12-14, Alg. 2].
Exploration Phase: If no feasible solution is found during the exploitation phase, the planner switches to exploration [line 10, Algo. 2]. In this mode, any exploration algorithm can be employed to generate paths $\mathbf{p}$} that optimize information-theoretic objectives (e.g., mutual information) subject to the constraints in (6); see, e.g., [19, 41]. Given such a path $\mathbf{p}$}, the planner selects ${\mathbf{x}}_{t+1}={\mathbf{p}}(1)$, sends it to the robot, updates the time step to $t+1$, and repeats the process [lines 9-14, Alg. 2]. The goal is for exploration paths to reduce uncertainty (i.e., the size of the prediction sets) sufficiently for the robot to re-enter exploitation mode and eventually complete $\psi$.
During both exploration and exploitation we also ensure that $\lVert{\mathbf{x}}_{t+1}-{\mathbf{x}}_{t}\rVert<r_{\text{max}}$ to eliminate cases where the next state may lie in the unexplored part of the map. The above process terminates once ${\mathbf{x}}_{t+1}\in{\mathcal{X}}_{g}$. Concatenation of these states across time yields ${\mathbf{x}}_{0:H}$.
@@
3:Collect calibration mission scenarios $\{\xi_{i}\}_{i=1}^{D}$
@@ IV Performance Guarantees
In this section, we show that our planner addresses Problem II.1, i.e., the path ${\mathbf{x}}_{0:H}$ generated by Algorithm 2 completes $\psi$ with probability $1-\alpha$. This result builds upon the fact that the sets ${\mathcal{C}}_{j}({\mathbf{m}}_{t})$, constructed by Alg. 1, satisfy (3), i.e., they jointly contain the true labels for all $j\in{\mathcal{J}}_{t}$, $t\in[0,H]$ and any path ${\mathbf{x}}_{0:H}$ with probability at least $1-\alpha$. This property is shown in the following proposition.
@@
Consider a test scenario $\xi_{\text{test}}=\{{\mathbf{x}}_{0,test},\Omega_{\text{test}},\psi_{\text{test}}\}\sim{\mathcal{D}}$ with unknown true labels $\{k_{j}^{\text{gt}}\}_{j\in{\mathcal{J}}}$. Given any path ${\mathbf{x}}_{0:H}$ in $\Omega_{\text{test}}$ yielding a sequence of maps ${\mathbf{m}}_{0:H}$, the prediction sets $\{{\mathcal{C}}_{j}({\mathbf{m}}_{t})\}_{j\in{\mathcal{J}}_{t}}$, computed at every $t\in\{0,\ldots,H\}$ constructed by Alg. 1 satisfy (3), for a user-defined $\alpha\in(0,1)$.
@@
Using a standard CP argument [30, 42], we have:
$\mathbb{P}_{\xi_{\text{test}}\sim{\mathcal{D}}}\left(s_{\text{test}}\leq\hat{s}\right|{\mathbf{m}}_{0:H})\geq 1-\alpha,$ (7)
where $s_{\text{test}}$ is the NCS associated with $\xi_{\text{test}}$. Note that $s_{\text{test}}$ is unknown as the ground truth labels $k_{j}^{\text{gt}}$ for cells $j$ are unknown. By definition of the NCS in (4), we have that $s_{\text{test}}=\max_{{\mathbf{x}}_{0:H}\in{\mathcal{P}},j\in{\mathcal{J}}_{t},t\in[0:H]}1-p_{t}(m_{j}=k_{j}^{gt})$. Plugging this into (7) yields the following result:
$\mathbb{P}_{\xi_{\text{test}}\sim{\mathcal{D}}}$ $\left(\max_{{\mathbf{x}}_{0:H}\in{\mathcal{P}},j\in{\mathcal{J}}_{t},t\in[0:H]}(1-p_{t}(m_{j}=k_{j}^{gt})\leq\hat{s}\mid{\mathbf{m}}_{0:H})\right)$
@@
$\mathbb{P}_{\xi_{\text{test}}\sim{\mathcal{D}}}$ $\biggl(\bigwedge_{\forall j\in{\mathcal{J}}_{t},\forall t\in[0:H],}p_{t}(m_{j}=k_{j}^{gt})\geq 1-\hat{s}\mid{\mathbf{m}}_{0:H}\biggr)$
@@
Recall from (5) that the prediction sets ${\mathcal{C}}_{j}({\mathbf{m}}_{t})$ collect all labels $k\in{\mathcal{K}}$ satisfying $p_{t}(m_{j}=k)\geq 1-\hat{s}$. Combining this with (IV) results in (3), i.e., $\mathbb{P}_{\xi_{\text{test}}\sim{\mathcal{D}}}\left(C_{0:H}|{\mathbf{m}}_{0:H}\right)\geq 1-\alpha$, where $C_{0:H}=\bigwedge_{\forall j\in{\mathcal{J}}_{t},t\in[0:H]}(k_{j}^{\text{gt}}\in{\mathcal{C}}_{j}({\mathbf{m}}_{t}))$. ∎
Building on Proposition IV.1, we show that the path ${\mathbf{x}}_{0:H}$ generated by Alg. 2 addresses Problem II.1. To show this result, we make the following assumptions about the exploration and planning algorithms employed in Alg. 2.
@@
(i) The exploration algorithm can reduce mapping uncertainty—and, consequently, the size of the prediction sets—so that the planner eventually enters the exploitation mode. This implies that exploration is invoked only a finite number of times before a path ${\mathbf{x}}_{0:H}$ is computed. This assumption guarantees that Alg. 2 eventually terminates with a path for any mission scenario $\xi\sim{\mathcal{D}}$ and, therefore, mission failures cannot be attributed to inefficient exploration. (ii) The planning algorithm used in the exploitation phase is complete, i.e., if a feasible path $\mathbf{p}$} exists that completes $\psi$ under the assigned labels described in Section III-B, then the planner is guaranteed to find it. This assumption ensures that mission failures can only be attributed to mapping errors.
@@
Given a test scenario $\xi_{\text{test}}=\{{\mathbf{x}}_{0,test},\Omega_{\text{test}},\psi_{\text{test}}\}\sim{\mathcal{D}}$ with unknown true labels $\{k_{j}^{\text{gt}}\}_{j\in{\mathcal{J}}}$, Alg. 2 will generate a path ${\mathbf{x}}_{0:H}$ satisfying (2).
@@
First, by Assumption IV.2, for any scenario $\xi_{\text{test}}\sim{\mathcal{D}}$, Algorithm 2 is guaranteed to terminate and return a path ${\mathbf{x}}_{0:H}$. Consequently, any failure to complete $\psi$ can be attributed only to mapping errors.
Second, consider the path ${\mathbf{x}}_{0:H}$ returned by Algorithm 2 for a scenario $\xi_{\text{test}}\sim{\mathcal{D}}$. We show that ${\mathbf{x}}_{0:H}$ satisfies (2) by leveraging Proposition IV.1. Specifically, by construction of the exploration and exploitation modes in Algorithm 2, Assumption IV.2-(ii), and Assumption II.2 we have that: (i) all states ${\mathbf{x}}_{t}$ in ${\mathbf{x}}_{0:H}$ satisfy the imposed safety constraints given the labels $k_{j}=\operatornamewithlimits{argmax}_{k\in{\mathcal{C}}_{j}({\mathbf{m}}_{t})}d_{k}$ assigned to cells $j$, i.e., $\max_{k\in{\mathcal{C}}_{j}({\mathbf{m}}_{t})}d_{k}+r_{m}\leq\lVert{\mathbf{x}}_{t+1}|_{\Omega}-c_{j}\rVert,\forall j\in{\mathcal{J}}_{t}$ and (ii) ${\mathbf{x}}_{H}\in{\mathcal{X}}_{g}$. Now, let us assume that the ground truth label is included in the prediction sets for all $t\in\{0,\ldots,H\}$. This equivalently means that the probability that ${\mathbf{x}}_{0:H}$ satisfies $\psi$ with respect to the assigned labels, and implicitly the ground truth labels, is $1$. The implicit satisfaction with respect to the ground truth is due to the consideration of the worst-case label in the prediction set that also contains the ground truth, In math, we express this as follows:
$\mathbb{P}_{\xi_{\text{test}}\sim{\mathcal{D}}}\left({\mathbf{x}}_{0:H}\models\psi|C_{0:H},{\mathbf{m}}_{0:H}\right)=1,$ (9)
where $C_{0:H}=\bigwedge_{\forall j\in{\mathcal{J}}_{t},t\in[0:H]}\left(k_{j}^{gt}\in{\mathcal{C}}_{j}({\mathbf{m}}_{t})\right)$. Recall from Proposition IV.1, that $\mathbb{P}_{\xi_{\text{test}}\sim{\mathcal{D}}}\left(C_{0:H}|{\mathbf{m}}_{0:H}\right)\geq 1-\alpha$. Combining this with (9) we get that ${\mathbf{x}}_{0:H}$ satisfies (2). ∎
@@
The guarantees in Proposition IV.1 and Theorem IV.3 are marginal, where the probabilities are over the randomness of the calibration set and the test scenario. A dataset-conditional guarantee which holds for a fixed calibration set across test scenarios can also be applied [43]. In such cases, Proposition IV.1 is modified such that with a probability of atleast $1-\delta$ over the sampling of the calibration set $\{\xi_{i}\}_{i=1}^{D}\sim{\mathcal{D}}$, $\mathbb{P}_{\xi_{\text{test}}\sim{\mathcal{D}}}\left({\mathcal{C}}_{0:H}|{\mathbf{m}}_{0:H},\{\xi_{i}\}_{i=1}^{D}\right)\geq Beta^{-1}_{D+1-v,v}(\delta),$ where $Beta^{-1}_{D+1-v,v}(\delta)$ is the $\delta$-quantile of the Beta distribution with parameters, $D$ - the size of the calibration dataset, and $v=\lfloor(D+1)\hat{\alpha}\rfloor$ and $\hat{\alpha}$ is selected so that $Beta^{-1}_{D+1-v,v}(\delta)\approx 1-\alpha$. Similarly, Theorem IV.3 is modified, where with a probability of at least $1-\delta$ over the sampling of the calibration dataset, we have that $\mathbb{P}_{\xi_{\text{test}}\sim{\mathcal{D}}}\left({\mathbf{x}}_{0:H}\models\psi|{\mathbf{m}}_{0:H},\{\xi_{i}\}_{i=1}^{D}\right)\geq Beta^{-1}_{D+1-v,v}(\delta)$.
@@ V Experimental Validation
In this section, we present experiments to validate the proposed approach. Section V-A defines the robot model and the distribution $\mathcal{D}$} used in our evaluation. Section V-B describes the configuration of the proposed algorithm and the evaluation metrics. The empirical performance of our method is presented in Section V-C, while Section V-D provides comparative results against uncertainty-informed and uncertainty-agnostic planners, showing that our approach outperforms the baselines in terms of mission completion rates. Finally, Section V-E reports success rates on out-of-distribution scenarios, highlighting potential failure cases. A video with an explanation of the framework, experiments and demonstrations can be found in [44].
V-A Setting Up Experiments: Robot & Distribution $\mathcal{D}$}
Robot: Our experiments are conducted using ROS 2 Humble [45] and the Gazebo (Ignition) simulator. The simulated platform is a Clearpath Husky skid-steer ground vehicle with a maximum linear speed of $1.0~\text{m/s}$ and a maximum yaw rate of $0.5~\text{rad/s}$. The robot is controlled at $2$ Hz. It is equipped with a forward-facing RGB-D camera with a resolution of $640\times 480$, a frame rate of $30$ Hz, a horizontal FoV of $90^{\circ}$, a vertical FoV of $73.7^{\circ}$, and a sensing range of $r_{\max}=10$ m. Note that this sensor is not omnidirectional, as required in Assumption II.2, due to its FoV; consequently, our algorithm is tested in a more challenging setting.

### v2

@@ III-B Planning over Conformalized Semantic Maps
Consider an environment $\Omega\sim\mathcal{D}$. Our goal is to compute a plan ${\mathbf{x}}_{0:H}$ satisfying any user-defined mission specification $\psi$ with probability at least $1-\alpha$. This path will be constructed incrementally as the map ${\mathbf{m}}_{t}$ is updated online; see Alg. 2. First, we collect the calibration dataset $\{\Omega_{i}\}_{i=1}^{D}$, and compute $\hat{s}$ as described in Section III-A [lines 3-4, Alg. 2]. Suppose at time $t>0$, the robot is at state ${\mathbf{x}}_{t}$, collects measurements ${\mathcal{Z}}_{t}$, and constructs the map ${\mathbf{m}}_{t}$ [lines 6-7, Alg. 2]. Using ${\mathbf{m}}_{t}$, the prediction sets ${\mathcal{C}}_{j}({\mathbf{m}}_{t})$ are constructed for all cells $j\in{\mathcal{J}}_{t}$ [line 8, Alg. 2]. We now discuss how the next robot state ${\mathbf{x}}_{t+1}$ is computed using these sets [lines 914, Alg. 2]. The next state ${\mathbf{x}}_{t+1}$ must be chosen so that it satisfies (1). Since the regions ${\mathcal{L}}_{k}$ are unknown, satisfaction of (1) must be checked using the map ${\mathbf{m}}_{t}$. Specifically, given ${\mathbf{m}}_{t}$, (1) can be re-written as:
$\lVert{\mathbf{x}}_{t+1}|_{\Omega}-c_{j}\rVert\geq d_{k_{j}}+r_{m},$ (6)
@@
Exploitation Phase: Using an existing motion planner [a-star, 12, 10], we first check whether there exists a robot path ${\mathbf{p}}={\mathbf{p}}_{0},{\mathbf{p}}_{1},\dots,{\mathbf{p}}_{T}$, with ${\mathbf{p}}_{n}\in{\mathcal{X}}$ for all $n\in\{0,\dots,T\}$ and some $T\geq 0$ that completes $\psi$ under the assigned labels $k_{j}$ [line 9, Alg. 2]. The path must satisfy ${\mathbf{p}}_{0}={\mathbf{x}}_{t}$, ${\mathbf{p}}_{T}\in{\mathcal{X}}_{g}$, and $\lVert{\mathbf{p}}_{n}|_{\Omega}-c_{j}\rVert\geq d_{k_{j}}+r_{m},\forall n$. If such a feasible plan exists, the robot selects ${\mathbf{x}}_{t+1}={\mathbf{p}}_{1}$, moves towards it, updates the time step to $t+1$, and repeats the process [lines 12-14, Alg. 2].
Exploration Phase: If no feasible solution is found during the exploitation phase, the planner switches to exploration [line 10, Algo. 2]. In this mode, any exploration algorithm can be employed to generate paths $\mathbf{p}$} that optimize information-theoretic objectives (e.g., mutual information) subject to the constraints in (6); see, e.g., [ssmi_temp, 4]. Given such a path $\mathbf{p}$}, the planner selects ${\mathbf{x}}_{t+1}={\mathbf{p}}_{1}$, sends it to the robot, updates the time step to $t+1$, and repeats the process [lines 9-14, Alg. 2]. The goal is for exploration paths to reduce uncertainty (i.e., prediction set sizes) for the robot to re-enter exploitation mode.
During both exploration and exploitation we also ensure that $\lVert{\mathbf{x}}_{t+1}-{\mathbf{x}}_{t}\rVert<r_{\text{max}}$ to eliminate cases where the next state may lie in the unexplored part of the map. The above process terminates once ${\mathbf{x}}_{t+1}\in{\mathcal{X}}_{g}$, and the robot reaches ${\mathbf{x}}_{t+1}$. Concatenation of these states across time yields ${\mathbf{x}}_{0:H}$.
@@
3:Collect calibration environments $\{\Omega_{i}\}_{i=1}^{D}$
@@ IV Performance Guarantees
In this section, we show that our planner addresses Problem II.1, i.e., the path ${\mathbf{x}}_{0:H}$ generated by Algorithm 2 completes $\psi$ with probability $1-\alpha$. This result builds upon the fact that the sets ${\mathcal{C}}_{j}({\mathbf{m}}_{t})$, constructed by Alg. 1, satisfy (3), i.e., they jointly contain the true labels for all $j\in{\mathcal{J}}_{t}$, $t\in[0,H]$ and any path ${\mathbf{x}}_{0:H}$ with probability at least $1-\alpha$.
@@
Consider a test environment $\Omega_{\text{test}}\sim{\mathcal{D}}$ with unknown true labels $\{k_{j}^{\text{gt}}\}_{j\in{\mathcal{J}}}$. Given any path ${\mathbf{x}}_{0:H}$ in $\Omega_{\text{test}}$ yielding a sequence of maps ${\mathbf{m}}_{0:H}$, the prediction sets $\{{\mathcal{C}}_{j}({\mathbf{m}}_{t})\}_{j\in{\mathcal{J}}_{t}}$, computed at every $t\in\{0,\ldots,H\}$ constructed by Alg. 1 satisfy (3), for a user-defined $\alpha\in(0,1)$.
@@
Using a standard CP argument [PwC, 1], we have:
$\mathbb{P}_{{\Omega_{\text{test}}\sim{\mathcal{D}}}}\left(s_{\text{test}}\leq\hat{s}\right|{\mathbf{m}}_{0:H})\geq 1-\alpha,$ (7)
where $s_{\text{test}}$ is the NCS associated with $\Omega_{\text{test}}$. Note that $s_{\text{test}}$ is unknown as the ground truth labels $k_{j}^{\text{gt}}$ for cells $j$ are unknown. By definition of the NCS in (4), we have that $s_{\text{test}}=\max_{{\mathbf{x}}_{0:H}\in{\mathcal{P}},j\in{\mathcal{J}}_{t},t\in[0:H]}1-p_{t}(m_{j}=k_{j}^{gt})$. Plugging this into (7) yields the following result:
$\mathbb{P}_{{\Omega_{\text{test}}\sim{\mathcal{D}}}}$ $\left(\max_{{\mathbf{x}}_{0:H}\in{\mathcal{P}},j\in{\mathcal{J}}_{t},t\in[0:H]}(1-p_{t}(m_{j}=k_{j}^{gt})\leq\hat{s}\mid{\mathbf{m}}_{0:H})\right)$
@@
$\mathbb{P}_{{\Omega_{\text{test}}\sim{\mathcal{D}}}}$ $\biggl(\bigwedge_{\forall j\in{\mathcal{J}}_{t},\forall t\in[0:H],}p_{t}(m_{j}=k_{j}^{gt})\geq 1-\hat{s}\mid{\mathbf{m}}_{0:H}\biggr)$
@@
Recall from (5) that the prediction sets ${\mathcal{C}}_{j}({\mathbf{m}}_{t})$ collect all labels $k\in{\mathcal{K}}$ satisfying $p_{t}(m_{j}=k)\geq 1-\hat{s}$. Combining this with (IV) results in (3), i.e., $\mathbb{P}_{{\Omega_{\text{test}}\sim{\mathcal{D}}}}\left(C_{0:H}|{\mathbf{m}}_{0:H}\right)\geq 1-\alpha$, where $C_{0:H}=\bigwedge_{\forall j\in{\mathcal{J}}_{t},t\in[0:H]}(k_{j}^{\text{gt}}\in{\mathcal{C}}_{j}({\mathbf{m}}_{t}))$. ∎
Observe that the result in Proposition IV.1 does not rely on Assumption II.2, as it focuses only on map cells $j\in\mathcal{J}_{t}$ for which at least one measurement has been taken. Building on Proposition IV.1, we show that the path ${\mathbf{x}}_{0:H}$ generated by Alg. 2 addresses Problem II.1. To show this result, we make the following assumptions about the exploration and planning algorithms employed in Alg. 2.
@@
(i) The exploration algorithm is invoked a finite number of times before a path ${\mathbf{x}}_{0:H}$ is computed. Each invocation reduces mapping uncertaity-and consequently prediction set size-so that the planner switches back to exploitation mode. (ii) The planning algorithm used in the exploitation phase is complete, i.e., if a feasible path $\mathbf{p}$} exists that completes $\psi$ under the assigned labels described in Section III-B, then the planner is guaranteed to find it. These assumptions ensure that mission failures can only be attributed to mapping errors.
@@
Consider a test environment $\Omega_{\text{test}}\sim{\mathcal{D}}$ with unknown true labels $\{k_{j}^{\text{gt}}\}_{j\in{\mathcal{J}}}$, and a mission specification $\psi_{\text{test}}$. Under Assumptions II.2 and IV.2, Alg. 2 will generate a path ${\mathbf{x}}_{0:H}$ satisfying (2).
@@
First, by Assumption IV.2, in any environment $\Omega_{\text{test}}\sim{\mathcal{D}}$, and for any task $\psi_{\text{test}}$, Algorithm 2 is guaranteed to terminate and return a path ${\mathbf{x}}_{0:H}$. Consequently, any failure to complete $\psi_{\text{test}}$ can be attributed only to mapping errors. Second, consider the path ${\mathbf{x}}_{0:H}$ returned by Algorithm 2 for the task $\psi_{\text{test}}$, when operating in $\Omega_{\text{test}}\sim{\mathcal{D}}$. We show that ${\mathbf{x}}_{0:H}$ satisfies (2) by leveraging Proposition IV.1. Specifically, by construction of the exploration and exploitation modes in Algorithm 2, Assumption IV.2-(ii), and Assumption II.2 we have that: (i) all states ${\mathbf{x}}_{t}$ in ${\mathbf{x}}_{0:H}$ satisfy the imposed safety constraints given the labels $k_{j}=\operatornamewithlimits{argmax}_{k\in{\mathcal{C}}_{j}({\mathbf{m}}_{t})}d_{k}$ assigned to cells $j$, i.e., $\max_{k\in{\mathcal{C}}_{j}({\mathbf{m}}_{t})}d_{k}+r_{m}\leq\lVert{\mathbf{x}}_{t+1}|_{\Omega}-c_{j}\rVert,\forall j\in{\mathcal{J}}_{t}$ and (ii) ${\mathbf{x}}_{H}\in{\mathcal{X}}_{g}$. Now, let us assume that the ground truth label is included in the prediction sets for all $t\in\{0,\ldots,H\}$. This equivalently means that the probability that ${\mathbf{x}}_{0:H}$ satisfies $\psi$ with respect to the assigned labels, and implicitly the ground truth labels, is $1$. The implicit satisfaction with respect to the ground truth is due to the consideration of the worst-case label in the prediction set that also contains the ground truth. We express this as follows:
$\mathbb{P}_{{\Omega_{\text{test}}\sim{\mathcal{D}}}}\left({\mathbf{x}}_{0:H}\models\psi|C_{0:H},{\mathbf{m}}_{0:H}\right)=1,$ (9)
where $C_{0:H}=\bigwedge_{\forall j\in{\mathcal{J}}_{t},t\in[0:H]}\left(k_{j}^{gt}\in{\mathcal{C}}_{j}({\mathbf{m}}_{t})\right)$. Recall from Proposition IV.1, that $\mathbb{P}_{{\Omega_{\text{test}}\sim{\mathcal{D}}}}\left(C_{0:H}|{\mathbf{m}}_{0:H}\right)\geq 1-\alpha$. Combining this with (9) we get that ${\mathbf{x}}_{0:H}$ satisfies (2). ∎
@@
[…]
@@ V Experimental Validation
In this section, we empirically validate the proposed approach. Sections V-A introduces the setup of our experiments. The empirical performance of our method is presented in Sections V-B-V-C, while Section V-D provides comparative results against uncertainty-informed and uncertainty-agnostic planners. Finally, Sections V-E-V-F evaluate our algorithm in out-of-distribution (OOD) real-world and simulation scenarios, highlighting potential failure cases. Demonstrations in simulation and real-world are provided in [6].
V-A Setting Up Experiments and the Proposed Planner
Robot: Our experiments are conducted using ROS 2 Humble [ros2] and the Gazebo (Ignition) simulator. The simulated platform is a Clearpath Husky skid-steer ground vehicle with a maximum linear speed of $1.0~\text{m/s}$ and a maximum yaw rate of $0.5~\text{rad/s}$. It is equipped with a forward-facing RGB-D camera with a resolution of $640\times 480$, a horizontal FoV of $90^{\circ}$, a vertical FoV of $73.7^{\circ}$, and a sensing range of $r_{\max}=10$ m. Note that this sensor is not omnidirectional, as required in Assumption II.2, due to its FoV; consequently, our algorithm is tested in a more challenging setting.